OpenStudy (anonymous):

Evaluate this exponential expression. (0.125)^2/3

OpenStudy (anonymous):

it easy

OpenStudy (jdoe0001):

what would be 0.125 as a fraction?

OpenStudy (jdoe0001):

so... let us convert to a fraction first let's see what we get

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

@jdoe0001 Now what?

OpenStudy (jdoe0001):

ohhh haemm ok one sec

OpenStudy (anonymous):

Answer choices: 0.0025 0.025 0.25

OpenStudy (jdoe0001):

\(\large { (0.125)^{\frac{2}{3}}\implies \left( \cfrac{1}{8} \right)^{\frac{2}{3}} \\ \quad \\ {\color{brown}{ 8=2^3}}\quad and\quad a^{-{\color{red} n}} \implies \cfrac{1}{a^{\color{red} n}}\qquad \qquad \cfrac{1}{a^{\color{red} n}}\implies a^{-{\color{red} n}}\quad thus \\ \quad \\ \left( \cfrac{1}{8} \right)^{\frac{2}{3}}\implies \left( \cfrac{1}{{\color{brown}{ 2^3}}} \right)^{\frac{2}{3}}\implies (2^{{\color{red}{ -3}}})^{\frac{2}{3}}\implies 2^{{\color{red}{ \cancel{-3}}}\frac{2}{\cancel{3}}}\implies ? }\)

OpenStudy (anonymous):

What?

OpenStudy (jdoe0001):

anything you don't follow there?

OpenStudy (anonymous):

Idk the answer. What is it? Lol

OpenStudy (jdoe0001):

well.. you tell us what is it?

OpenStudy (anonymous):

Here's the answer choices: 0.0025 0.025 0.25

OpenStudy (anonymous):

None of your work shows anything like the answer choices.

OpenStudy (jdoe0001):

so.. what does \(\large { (0.125)^{\frac{2}{3}}\implies \left( \cfrac{1}{8} \right)^{\frac{2}{3}} \\ \quad \\ {\color{brown}{ 8=2^3}}\quad and\quad a^{-{\color{red} n}} \implies \cfrac{1}{a^{\color{red} n}}\qquad \qquad \cfrac{1}{a^{\color{red} n}}\implies a^{-{\color{red} n}}\quad thus \\ \quad \\ \left( \cfrac{1}{8} \right)^{\frac{2}{3}}\implies \left( \cfrac{1}{{\color{brown}{ 2^3}}} \right)^{\frac{2}{3}}\implies (2^{{\color{red}{ -3}}})^{\frac{2}{3}}\implies 2^{{\color{red}{ \cancel{-3}}}\cdot \frac{2}{\cancel{3}}}\implies ? }\) give you?

OpenStudy (anonymous):

I have no idea

OpenStudy (jdoe0001):

the choices are not meant to be a menu to choose from they're just a guideline so... simplify that, see what you get, and the choices must reflect that

OpenStudy (anonymous):

Not really. It has to be one of them.

OpenStudy (anonymous):

Idk what you're meaning by any of that.

OpenStudy (jdoe0001):

well... you're expected to have studied all that by now, otherwise, makes no sense to give you the exercise

OpenStudy (anonymous):

I do online school so yeah. I'm just now learning it and the way you're explaining it doesn't make sense.

OpenStudy (jdoe0001):

unless you're expected to use the calculator and you may well could do that I gather

OpenStudy (anonymous):

2*2=4. That's what I got from your drawing.

OpenStudy (anonymous):

Is 4 one of the answer choices? Nope.

OpenStudy (jdoe0001):

2*2? hmmm well..... ok.... do you see how we got the \(2^3\) and then the \(2^{-3} ?\)

OpenStudy (anonymous):

No.

OpenStudy (jdoe0001):

ok how about this part \(\large {\color{brown}{ 8=2^3}}\quad and\quad a^{-{\color{red} n}} \implies \cfrac{1}{a^{\color{red} n}}\qquad \qquad \cfrac{1}{a^{\color{red} n}}\implies a^{-{\color{red} n}} \) anything confusing there?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

You're not using any of the numbers in MY equation. Lol

OpenStudy (jdoe0001):

ok... what part? is just notation, I understand hehe

OpenStudy (anonymous):

Any of it. Can you lead me to the answer or no?

OpenStudy (jdoe0001):

well... the idea being \(\large and\quad a^{-{\color{red} n}} \implies \cfrac{1}{a^{\color{red} n}}\qquad \qquad \cfrac{1}{a^{\color{red} n}}\implies a^{-{\color{red} n}}\) do you follow what that means? means, a number, with a negative exponent, could also be written as the same number with a positive exponent but under 1 or the other way around so any number with whatever exponent, could be rewritten as the same number, as denominator, with the exponent as negative

OpenStudy (anonymous):

But what about MY problem?

OpenStudy (jdoe0001):

dohh.. . rather what's \(\large 2^3=?\)

OpenStudy (anonymous):

8

OpenStudy (jdoe0001):

so... we know that \(\large 8= 2^3\) and we also know that we could write 1/8 as \(\cfract{1}{2^3}\) and we also know we could write \(\bf \cfrac{1}{2^3}\implies 2^{-3}\) so \(\bf (0.125)^{\frac{2}{3}}\implies \left( \cfrac{1}{8} \right)^{\frac{2}{3}} \\ \quad \\ {\color{brown}{ 8=2^3}}\quad and\quad a^{-{\color{red} n}} \implies \cfrac{1}{a^{\color{red} n}}\qquad \qquad \cfrac{1}{a^{\color{red} n}}\implies a^{-{\color{red} n}}\quad thus \\ \quad \\ \left( \cfrac{1}{8} \right)^{\frac{2}{3}}\implies \left( \cfrac{1}{{\color{brown}{ 2^3}}} \right)^{\frac{2}{3}}\implies (2^{{\color{red}{ -3}}})^{\frac{2}{3}}\) follow that?

OpenStudy (anonymous):

Nope

OpenStudy (jdoe0001):

hmmm so.. how are you expected to evaluate it then? I gather you haven't covered this section yet of negative exponentials so. what section is this exercise meant to cover? are you simply supposed to use the calculator? because you could do that too

OpenStudy (jdoe0001):

well...then you don't need help with it at least not as most folks here define it

OpenStudy (anonymous):

I do need help. You're just not very good at explaining it, obviously.

OpenStudy (anonymous):

Since I still don't understand it.